OpenStudy (tjonessss77):

You put $300 at the end of each month in an investment plan that pays an APR of 7%. How much will you have after 18 years? Compare this amount to the total deposits made over the time period. a. $129,201.10; $64,800 c. $129,216.31; $64,800 b. $129,211.25; $64,775 d. $129,218.51; $64,775

jimthompson5910 (jim_thompson5910):

This is a future value of annuity problem http://www.investopedia.com/terms/f/future-value-annuity.asp \[\Large FV = C*\left[\frac{(1+i)^n-1}{i}\right]\] FV = future value after n periods C = cashflow (aka amount of money deposited each time) i = interest rate per period n = number of periods In this case, FV = unknown C = 300 i = 7%/12 = 0.07/12 = 0.00583333333333 n = 18*12 = 216 Now plug this into the formula given on the link. Then use a calculator to compute \[\Large FV = C*\left[\frac{(1+i)^n-1}{i}\right]\] \[\Large FV = 300*\left[\frac{(1+0.00583333333333)^{216}-1}{0.00583333333333}\right]\] \[\Large FV \approx 129,216.307980597\] \[\Large FV \approx 129,216.31\] Don't forget to round to the nearest penny

jimthompson5910 (jim_thompson5910):

So after 18 years, you'll have $129,216.31 in the account This is assuming you deposit $300 each month (every month for 18 years). Compare this with just storing $300 under your mattress and you would have 300*12*18 = $64,800

jimthompson5910 (jim_thompson5910):

So clearly depositing the $300 into a bank account is the better choice. You practically double your money.